# A Comprehensive Survey on Vector Database: Storage and Retrieval Technique, Challenge

### v2

@@ V-F Data Security and Privacy Protection
As cyber threats intensify and regulatory requirements become more stringent, data security and privacy protection for databases have become top priorities[103]. Compared to the comprehensive data security and privacy protection features of traditional relational databases, VDBs are still in the early stages. Unlike traditional relational databases, VDBs typically handle large amounts of high-dimensional embedding vectors, which pose higher risks of privacy breaches during storage, querying, and transmission. This is especially true in cloud platforms, where data is often stored in an unencrypted form or transmitted between different nodes, increasing the potential attack surface[104, 105]. In the future, VDBs will need to not only leverage traditional data security and privacy protection methods to build a robust security framework, but also integrate emerging technologies, such as blockchain tables[106] (to ensure data immutability) and AI-based anomaly detection (for proactive threat management), to adapt to specific application scenarios.
VI Synergy of LLMs and VDBs
By virtue of its excellent capability for rapid processing of unstructured data, a VDB can naturally meet the requirements of vector-intensive applications, especially for LLMs, where its role is becoming increasingly crucial. When processing natural language, LLMs need to convert text into high-dimensional vectors for computation and analysis, which demands robust storage and retrieval capabilities for high-dimensional vectors. Meanwhile, alleviating issues such as hallucinations and forgetfulness in LLMs also requires supporting facilities like vast external knowledge bases, all of which necessitate the support of VDBs. Therefore, the integration of LLMs and VDBs is set to be an inevitable trend in the future. Consequently, this section will delve into the integration and mutual influence between VDBs and LLMs, and provide an outlook on relevant potential applications, with the aim of offering valuable references and insights for subsequent scientific research and industrial applications. In the following sections, We will delve into the synergistic interaction between VDBs and LLMs, exploring their concrete application prospects in depth.
Given these challenges, researchers and developers have been exploring innovative solutions to enhance the performance and reliability of LLMs. One promising approach is the integration of VDBs into LLM systems.
VI-A VDBs for LLMs
@@
LLMs are characterized by large model capacity and vast data corpus[107]. With hundreds of billions (or more) of parameters and extensive textual training, they are highly adept at comprehending human knowledge and instructions[108]. However, LLMs do have certain shortcomings, though[3]. One major shortcoming is hallucinations where the model generates a response that is factually inaccurate. This shortcoming is mainly caused by the following issues, including knowledge limitations confined by the training corpa, the internal knowledge in LLMs cannot be updated resulting in outdated knowledge, LLMs may also introduce systematic errors due to the large dataset used for training. Another shortcoming is oblivion problem. LLMs have been found having the inclination to forget the previous input information, and also exhibit catastrophic forgetting behavior. In response to these issues, VDBs can offer robust support for LLMs in the following aspects:
1) VDBs as an External Knowledge Base: Retrieval-Augmented Generation (RAG). The Retrieval-Augmented Generation (RAG) technique is an artificial intelligence technology that combines information retrieval technology with language generation models[109]. This technique enhances the capability of LLMs in handling knowledge-intensive tasks, such as question answering, text summarization, and content generation, by retrieving relevant information from an external knowledge base and inputting it as a prompt to the LLMs. To address the above limitations, recent research introduced the RAG technique. Retrieval models play an important role in various knowledge-intensive tasks by providing timely and accurate external knowledge through effective data maintenance in external databases. RAG models provide a promising direction for enhancing the performance and adaptability of LLMs in a variety of tasks by integrating retrieval with generation.
The RAG framework has become a paradigm and has brought a huge shift to NLP[110]. RAG models consist of several major processes in the era of LLMs, including retrieval, generation, and augmentation. A common workflow of RAG when meeting LLMs is illustrated in figure8. The complete operational workflow of the system essentially consists of three core components: data storage, information retrieval, and content generation.
@@
The RAG workflow begins with the data storage phase. During this phase, externally collected unstructured data (text, images, audio, video) undergoes preprocessing. The processed data is then divided into smaller chunks, converted into vectors via an embedding model to capture semantic representations, and stored in a VDB for subsequent vector retrieval.
@@
Next is the information retrieval phase. This stage starts when a user poses a question to the model in the form of a prompt. The embedding model (used earlier for processing external unstructured data) generates an embedding vector for the query, which is then used to retrieve the most semantically similar data chunks from the VDB. These retrieved results are converted back from vector format to their original format and returned to the user.
@@
Finally, in the content generation phase, the large language model (LLM) generates the final answer. The user’s original question and the retrieved information are integrated into a task-specific prompt template (the selection of the prompt template depends on the task type). The LLM then processes this prompt and produces the answer.
@@
VDBs as a Cost-effective Semantic Cache. The running of LLMs consume huge resources. Its training requires massive computing power,and frequent API calls to third-party models racks up significant costs. With the help of VDBs, the interaction cost and computing workload of LLMs can be reduced significantly, thereby promoting cost-effective and end-to-end applications of LLMs.
By integrating VDBs with LLMs, VDBs serve as GPT semantic cache, which leverages semantic caching of query embeddings in in-memory storage. This method can efficiently identify semantically similar questions by storing embeddings of user queries, allowing for the retrieval of pre-generated responses without redundant API calls to the LLMs. This technique is an efficient way to reduce operational costs ,improve response times and address inefficiency[111]. This architecture mainly consists of three components, including an embedding generation that converts user queries into semantic embeddings, an in-memory caching that manages storage and retrieval of embeddings and responses and a similarity search that identifies semantically similar queries. The combination of VDBs with LLMs offers several advantages. This technique reduces API dependency while maintaining high response accuracy. In addition, it demonstrates substantial scalability and adaptability. By using vector searching algorithms and in-memory caching, it allows for the handling of large volumes of queries without a proportional increase in computational demands, supporting stable performance even under fluctuating workloads. It also supports for multiple embedding models and configurations, making the system adaptable to various deployment needs[112]. Using VDBs as GPT semantic cache will be a viable solution to facilitate the large-scale application of LLMs.
VDBs as A Reliable Memory of LLMs. Memory systems can power the intelligence LLMs, enabling them to demonstrate the capability of autonomous and thus show impressive performance in a wide range of tasks. The integration of memory systems and LLMs is conductive to the coherence, contextual, and efficiency of interactions and that the system can learn and adapt over time. Currently, a significant drawback of LLMs is lacking strong long-term memory capabilities[113]. This limitation will hinder LLMs’ ability to maintain context over long periods of time and retrieve relevant information from past interactions. Therefore, in order to improve the decision quality and reasoning efficiency of LLMs in complex tasks, it is necessary to research and develop effective long-term memory mechanisms. Through the external knowledge storage and historical interactions, long-term memory is available for LLMs to store and retrieve and use in subsequent interactions, which enhances LLMs’ intelligence ability in maintaining contextual coherence, improving decision-making quality, reducing cost of reasoning and demonstrating higher intelligence in long-term interactions.
The VDBs can be used as the underlying basic tool of LLMs to support the storage of historical information, so that LLMs can effectively store different types of historical interaction information, such as knowledge information, dialogue information, and related task information. Then different types of information during intelligent interaction are stored in the VDBs as long term memory after slicing and vectorizing, which facilitates the subsequent retrieval and updating of relevant memory information. Another shortcoming of LLMs is that they cannot update knowledge dynamically, lacking of the few-shot learning ability. VDBs provide a robust memory layer for LLMs to update new information continuously in the way of storage, thus ensuring that LLMs can make response according to the most current and relevant data.
VI-B LLMs for VDBs
@@
In addition, LLMs in turn can empower databases.AI technology has been proved to perform well in many data management tasks, such as data processing, database optimization, and data analysis. However, traditional machine learning algorithms are unable to solve generalization and inference problems. For example, traditional machine learning algorithms have difficulty in adapting to different databases, different query workloads, and different hardware environments, making them unable to solve the generalizability and inference problems in data management tasks. In addition, traditional machine learning algorithms cannot satisfy the need for contextual understanding and multi-step reasoning in optimization scenarios such as database diagnosis, root cause analysis, etc. However, LLMs bring promising solutions to the above problems[114].
LLMs assist database management tasks. LLMs revolutionize data management. LLMs show great potential in optimizing data management problems due to its excellent language comprehension and generalization capabilities in tasks such as data processing, database optimization, and data analysis. For example, LLMs can analyze anomalous database metrics and report root causes and potential solutions to data base administrators. LLMs can also be used as a natural language (NL) interface for data analysis tasks, converting NL requests into executable queries against their databases. LLMs bring several advantages to database management tasks. The first point is higher transfer capability. Existing instance-optimal works can optimize an instance but cannot be extended to other instances, whereas, the combination of LLMs with databases demonstrates exceptional transfer capability. With a few fine-tuned examples, comparable performance can be achieved on novel database tasks, making them more adaptable to database schema, workload, or even data and hardware changes. The second point is the ability to provide a user-friendly interface. LLMs allow users to provide some prompts as hints to guide the model’s inference, thus offering an intuitive interface. The second is the ability to provide a user-friendly interface. LLMs allow users to provide some prompts as hints to guide the model’s inference, thus offering an intuitive user experience without the need for large amounts of training data (supervised modelling) or multiple iterations (reinforcement learning) to capture and merge user feedback. The third point is to learn from prior knowledge. LLMs can extract insights from existing database components, including documents and even code. By integrating the strengths of these components, databases’ performance can be enhanced while mitigating the individual weaknesses of the components[115].
@@
LLMs smarten vector data handling. The deep integration of LLMs with VDBs has pioneered innovative application scenarios for data-driven workflows, encompassing content generation, knowledge enhancement, and system optimization. By combining semantic understanding with vectorized retrieval, LLMs can generate customized texts (e.g., thematic articles, stylized summaries) based on vector inputs, enrich ambiguous texts with additional details (e.g., supplementing statistical data or case studies), and facilitate cross-language, cross-domain text transformations (e.g., multilingual simplification of legal documents). Furthermore, LLMs significantly optimize the management tasks of VDBs: they recommend configuration parameters by analyzing historical performance data to improve system stability, automatically diagnose performance bottlenecks while generating interpretable reports, and efficiently process heterogeneous data through semantic analysis (e.g., schema matching and error correction). These applications not only reduce the cost of manual intervention but also extend the generalization capabilities of traditional methods through adaptive solutions, highlighting the core value of LLMs in enhancing the intelligence and scalability of VDBs[116, 117, 118, 119, 120, 121, 122].
@@
VI-C A General LLMs and VDBs Synergized Framework
For a framework that incorporates a large language model and a VDB, as shown in figure8 it can be understood by splitting it into four levels: the user level, the model level, the AI database level, and the data level, respectively. For a user who has never been exposed to large language modeling, it is possible to enter natural language to describe their problem. For a user who is proficient in large language modeling, a well-designed prompt can be entered. The LLM next processes the problem to extract the key- words in it, or in the case of open source LLMs, the corre- sponding vector embeddings can be obtained directly. The VDB stores unstructured data and their joint embeddings. The next step is to go to the VDB to find similar nearest neighbors. The ones obtained from the sequences in the big language model are compared with the vector encodings in the VDB, by means of the NNS or ANNS algorithms. And different results are derived through a predefined serialization chain, which plays the role of a search engine. If it is not a generalized question, the results derived need to be further put into the domain model, for example, imagine we are seeking an intelligent scientific assistant, which can be put into the model of AI4S to get professional results. Eventually it can be placed again into the LLM to get coherent generated results. For the data layer located at the bottom, one can choose from a variety of file formats such as PDF, CSV, MD, DOC, PNG, SQL, etc., and its sources can be journals, conferences, textbooks, and so on. Corresponding disciplines can be art, science, engineering, business, medicine, law, and etc.
VII Conclusion
In this paper, we provide a comprehensive and up-to-date literature review on VDBs, including the key algorithms, storage, and retrieval methods. We also compare representative VDB systems, analyze their design trade-offs, and discuss their strengths, limitations, and typical use cases. Furthermore, we identify key challenges and outline potential research directions, including improved indexing and closer integration with LLMs. We believe this survey offers a solid reference for researchers and practitioners, and contributes to a clearer understanding of the current state and future direction of vector databases.

### v3

@@ 5.5 Data Security and Privacy Protection
As cyber threats intensify and regulatory requirements become more stringent, data security and privacy protection for databases have become top priorities [farayola2024data]. Compared to the comprehensive data security and privacy protection features of traditional relational databases, VDBs are still in the early stages. Unlike traditional relational databases, VDBs typically handle large amounts of high-dimensional embedding vectors, which pose higher risks of privacy breaches during storage, querying, and transmission. This is especially true in cloud platforms, where data is often stored in an unencrypted form or transmitted between different nodes, increasing the potential attack surface [asaad2024enhancing, amaithi2024systematic]. In the future, VDBs will need to not only leverage traditional data security and privacy protection methods to build a robust security framework but also integrate emerging technologies, such as blockchain tables [s23167172] (to ensure data immutability) and AI-based anomaly detection (for proactive threat management), to adapt to specific application scenarios.
6 Synergy of LLMs and VDBs
When processing natural language, Large Language Models (LLMs) need to convert natural language into high-dimensional vectors, which requires robust capabilities for storing and retrieving high-dimensional vectors. Meanwhile, the inherent issues of LLMs, such as "hallucinations" (generating content that seems plausible but is factually incorrect) and "forgetfulness" (difficulty in remembering or utilizing early information), also necessitate supporting facilities like vast external knowledge bases to mitigate them, all of which cannot be achieved without the support of VDBs. Therefore, the integration of LLMs and VDBs is an inevitable trend. Among the approaches, integrating VDBs into LLM systems is a promising method. In the following section, we will delve into the fusion, mutual influence, and potential applications of the two, providing references for scientific research and industrial applications.
6.1 VDBs for LLMs
@@
LLMs are characterized by large model capacity and vast data corpus[openai2024gpt]. With hundreds of billions (or more) of parameters and extensive textual training, they are highly adept at comprehending human knowledge and instructions [shanahan2023talking]. However, LLMs do have certain shortcomings, though [zhao2023survey]. One major shortcoming is hallucinations, where the model generates a response that is factually inaccurate. This shortcoming is mainly caused by the following issues, including knowledge limitations confined by the training corpus; the internal knowledge in LLMs cannot be updated, resulting in outdated knowledge; and LLMs may also introduce systematic errors due to the large dataset used for training. Another shortcoming is the oblivion problem. LLMs have been found to have the inclination to forget the previous input information and also exhibit catastrophic forgetting behavior. In response to these issues, VDBs can offer robust support for LLMs in the following aspects:
1) VDBs as an External Knowledge Base: Retrieval-Augmented Generation (RAG). VDBs as External Knowledge Base: Retrieval-Augmented Generation (RAG). Retrieval-Augmented Generation (RAG) technology is an innovation in the field of artificial intelligence that integrates information retrieval with language generation models [lewis2020retrieval, lewis2021retrievalaugmentedgenerationknowledgeintensivenlp]. This technology significantly enhances the performance of Large Language Models (LLMs) in knowledge-intensive tasks such as question answering, text summarization, and content generation by retrieving relevant information from an external knowledge base and inputting it as a prompt to LLMs. RAG encompasses core processes including retrieval, generation, and augmentation. Figure 8 illustrates a typical workflow of RAG when integrated with LLMs. The system’s operational workflow mainly consists of three core components: data storage, information retrieval, and content generation.
@@
The RAG workflow begins with the data storage phase. During this phase, externally collected unstructured data (text, images, audio, and video) undergoes preprocessing, is divided into smaller chunks, and is converted into vectors through an embedding model to capture semantic representations. These vectors are then stored in a VDB for subsequent vector retrieval.
@@
Next is the information retrieval phase. When a user poses a question in the form of a prompt, the embedding model generates a vector for the query and retrieves the most semantically similar data chunks from the VDB. The retrieved results are converted back from vector format to their original format and returned to the user.
@@
Finally, in the content generation phase, the Large Language Model (LLM) integrates the user’s original question and the retrieved information, selects an appropriate prompt template based on the task type, processes the prompt, and generates the final answer.
@@
2) VDBs as a Cost-effective Semantic Cache. The operation of LLMs requires substantial computational resources, and frequent API calls to third-party models incur high costs. By integrating VDBs, they can be utilized as a semantic cache to achieve efficient semantic matching through query embeddings. This method, by storing query embeddings, can reduce redundant API calls, lower costs, and improve response speed and efficiency [regmi2024gpt]. Moreover, the combination of VDBs and LLMs demonstrates strong scalability and adaptability, enabling efficient handling of a large volume of queries, stable performance under fluctuating workloads, and support for multiple embedding models and configurations to meet diverse deployment needs  [bang2023gptcache]. Using VDBs as a semantic cache for GPT is a feasible strategy to promote the large-scale application of LLMs.
3) VDBs as a Reliable Memory of LLMs. A notable shortcoming of current LLMs is their lack of strong long-term memory capabilities [hatalis2023memory]. Memory systems can enhance the intelligence of LLMs, enabling them to possess autonomous capabilities and improve their performance in various tasks. VDBs can act as a memory tool for LLMs, supporting the storage of historical information and allowing LLMs to effectively store different types of historical interaction information, such as knowledge, dialogue, and related task information. Additionally, LLMs lack the ability to dynamically update knowledge and engage in few-shot learning. VDBs can continuously update new information to ensure that responses are made based on the latest and most relevant data.
6.2 LLMs for VDBs
@@
In addition, LLMs in turn can empower databases. AI technology has been proved to perform well in many data management tasks, such as data processing, database optimization, and data analysis. However, traditional machine learning algorithms are unable to solve generalization and inference problems. For example, traditional machine learning algorithms have difficulty in adapting to different databases, different query workloads, and different hardware environments, making them unable to solve the generalizability and inference problems in data management tasks. In addition, traditional machine learning algorithms cannot satisfy the need for contextual understanding and multi-step reasoning in optimization scenarios such as database diagnosis, root cause analysis, etc. However, LLMs bring promising solutions to the above problems [zhou2024db].
LLMs assist database management tasks. Large Language Models (LLMs) are actively contributing to database management and driving innovation in the field of data management. For instance, LLMs can analyze abnormal database metrics, report the root causes and potential solutions to administrators, and also serve as a natural language interface to convert user requests into executable queries. Moreover, after being integrated with databases, LLMs can perform exceptionally well on new tasks with minimal fine-tuning, making them more adaptable to changes in database schemas, data, and hardware. By guiding model inference through prompts, LLMs achieve a user-friendly interface, providing an intuitive experience without the need for extensive training data or multiple iterations to incorporate user feedback. Additionally, LLMs can extract insights from database components (such as documents and code), integrating their strengths to enhance database performance and compensate for the weaknesses of individual components [li2024llm].
@@
LLMs smarten vector data handling. The deep integration of LLMs with VDBs has pioneered innovative application scenarios for data-driven workflows, encompassing content generation, knowledge enhancement, and system optimization. By combining semantic understanding with vectorized retrieval, LLMs can generate customized texts (e.g., thematic articles, stylized summaries) based on vector inputs, enrich ambiguous texts with additional details (e.g., supplementing statistical data or case studies), and facilitate cross-language, cross-domain text transformations (e.g., multilingual simplification of legal documents). Furthermore, LLMs significantly optimize the management tasks of VDBs: they recommend configuration parameters by analyzing historical performance data to improve system stability, automatically diagnose performance bottlenecks while generating interpretable reports, and efficiently process heterogeneous data through semantic analysis (e.g., schema matching and error correction). These applications not only reduce the cost of manual intervention but also extend the generalization capabilities of traditional methods through adaptive solutions, highlighting the core value of LLMs in enhancing the intelligence and scalability of VDBs [tang2023does, albalak2024survey, fan2024cost, zhou2023d, huang2024llmtune, chang2023prompt, whitehouse2023llmpowered].
@@
6.3 A General LLMs and VDBs Synergized Framework
For a framework that incorporates a large language model and a VDB, as shown in Figure 8, can be understood by splitting it into four levels: the user level, the model level, the AI database level, and the data level, respectively. For a user who has never been exposed to large language modeling, it is possible to enter natural language to describe their problem. For a user who is proficient in large language modeling, a well-designed prompt can be entered. The LLM next processes the problem to extract the key- words in it, or in the case of open-source LLMs, the corre- sponding vector embeddings can be obtained directly. The VDB stores unstructured data and their joint embeddings. The next step is to go to the VDB to find similar nearest neighbors. The ones obtained from the sequences in the big language model are compared with the vector encodings in the VDB by means of the NNS or ANNS algorithms. And different results are derived through a predefined serialization chain, which plays the role of a search engine. If it is not a generalized question, the results derived need to be further put into the domain model; for example, imagine we are seeking an intelligent scientific assistant, which can be put into the model of AI4S to get professional results. Eventually it can be placed again into the LLM to get coherent generated results. For the data layer located at the bottom, one can choose from a variety of file formats such as PDF, CSV, MD, DOC, PNG, SQL, etc., and its sources can be journals, conferences, textbooks, and so on. Corresponding disciplines can be art, science, engineering, business, medicine, law, etc.
7 Conclusion
In this paper, we provide a comprehensive and up-to-date literature review on VDBs, including the key algorithms, storage, and retrieval methods. We also compare representative VDB systems, analyze their design trade-offs, and discuss their strengths, limitations, and typical use cases. Furthermore, we identify key challenges and outline potential research directions, including improved indexing and closer integration with LLMs. We believe this survey offers a solid reference for researchers and practitioners, and contributes to a clearer understanding of the current state and future direction of VDBs.